OpenStudy (anonymous):

Which ordered pair is a solution of the equation 4x - 2y = 16? Select one: a. (-5, -3) b. (6, -4) c. (-5, 3) d. (6, 4)

OpenStudy (anonymous):

you basically have to substitute the x and y points given into the equation to see if it makes sense , e.g. (-5,-3) , x = -5 and y=-3 now lets substitute into the equation : 4x-2y = 4(-5) - 2(-3) = -20 -(-6) = -20+6 = -14 so this is does not give us 16, so we can rule out this option similarly , solve for the rest and you will get the answer

OpenStudy (anonymous):

so D?

OpenStudy (anonymous):

yup :)